# Pseudo-Simulation for Autonomous Driving

### v1

@@ 1 Introduction
However, accurate simulation remains a significant challenge. Real-world driving is visually complex and behaviorally diverse, making it difficult to replicate in simulation. Most existing platforms are manually constructed by 3D artists and engineers [3, 5]. This limits their realism and the diversity across scenes. Moreover, simulation-based evaluation is a computationally intensive and inherently sequential process. It often relies on large amounts of correlated evaluation frame sequences due to the high frequency of simulation required to ensure fidelity (usually 10Hz or higher).
Open-loop evaluation, on the other hand, assesses planning performance by comparing predicted trajectories to expert demonstrations in pre-recorded datasets. Each observation for evaluation includes sensor inputs, a goal location, and the future trajectory executed by a human expert driver. The AV predicts a fixed-horizon trajectory conditioned on the inputs, which is then scored against the expert using either displacement errors or metrics derived from ground-truth (GT) environment annotations, such as lane compliance or estimated collisions [6, 7, 8]. This approach operates entirely on real sensor data and avoids the complexities of interactive simulation, making it scalable and straightforward to apply over large datasets. However, it evaluates behavior only under expert-aligned conditions and does not account for distribution shifts. In deployment, the AV deviates from the demonstrated path, and open-loop protocols do not test its ability to recover from such drift.
@@
Using the nuPlan simulator [4], we demonstrate that pseudo-simulation achieves strong correlation with closed-loop results for a set of 83 diverse planners ($R^{2}=0.8$), while being substantially more efficient (6$\times$ less planner inferences and environment interactions). To enable standardized benchmarking, we release NAVSIM v2, a framework for benchmarking autonomous driving built upon our proposed evaluation methodology. We find that it reveals previously unknown failure modes in popular AV algorithms [11, 12], thus establishing it as a challenging new testbed for future research. We hope that pseudo-simulation can accelerate AV development through more efficient experimentation cycles and ensuring that future models prioritize robustness.
@@ 2 Related Work
Closed-Loop Benchmarking. Graphics-based simulators support closed-loop evaluation, but are computationally expensive and introduce domain gaps in sensor fidelity [3, 5]. To improve scalability, data-driven planning simulators leverage recorded traffic data [4, 24, 25, 26]. However, these systems operate at the trajectory level and do not support sensor-based agents. Several works attempt to bridge this gap through data-driven sensor simulation, generating synthetic views from real-world logs. Early systems simulate ego-vehicle deviations via image-based rendering [27, 28, 29]. More recent methods explore neural rendering [30, 31, 32, 33, 34, 35]. However, they face challenges related to photorealism, multi-agent consistency, and runtime efficiency. As such, there is no widely established neural rendering based AV benchmark yet, which we aim to address in this work.
Open-Loop Benchmarking. Open-loop evaluation typically measures planning quality via displacement errors between predicted and expert trajectories [36]. These metrics are simple to compute, but often correlate poorly with real-world performance and tend to favor trivial or history-based baselines [7, 12, 37, 38]. Furthermore, benchmarks adopting the nuScenes dataset [39] exhibit inconsistencies in metric implementation [6, 7] and overrepresent low-complexity scenes such as straight driving [7]. The closest work to ours, NAVSIM v1 [8], offers a more structured framework for open-loop benchmarking. The agent-under-test predicts a fixed-horizon trajectory from real sensor inputs, while other actors replay their recorded motion. This setup supports scalable evaluation and enables simulation-based metrics such as progress and collision rates. However, unlike pseudo-simulation, NAVSIM v1 remains limited to open-loop evaluation from expert-aligned initial observations and does not account for compounding errors or causal confusion [40].
@@ 3 Pseudo-Simulation
We now introduce the pseudo-simulation protocol for evaluating AV systems. Evaluation proceeds in two stages, as detailed in Section 3.1 and Section 3.2, and illustrated in Fig. 1.
Task Description. We consider a fixed-horizon planning task where in both Stage 1 and Stage 2, an AV (also called planner/ego agent) generates a 4-second trajectory based on sensor inputs and a driving command [8]. The inputs include multi-view camera images and ego status features such as the velocity and motion history. The driving command specifies the intended maneuver in case of ambiguity, e.g., at intersections, and is provided as a discrete label: left, straight, or right. The ego agent outputs a trajectory (i.e., a sequence of desired future waypoints) in its local coordinate frame.
@@ 3.1 Stage 1: Initial Observations
Here, $\mathcal{M}_{\text{pen}}=\{\text{NC},\text{DAC},\text{DDC},\text{TLC}\}$ and $\mathcal{M}_{\text{avg}}=\{\text{TTC},\text{EP},\text{HC},\text{LK},\text{EC}\}$ (Table 1). Unlike prior work [44], to prevent penalizing contextually justified maneuvers, we introduce a novel filtering mechanism for the EPDMS. If a rule violation is also committed by the human expert driver in the same scene, the penalty is ignored. This avoids penalizing infractions due to label noise or valid behaviors, such as briefly entering the opposite lane to bypass a static obstacle:
$\text{filter}_{m}(\text{agent},\text{human})=\begin{cases}1.0&\text{if }m(\text{human})=0\\ m(\text{agent})&\text{otherwise}\end{cases}$ (2)
@@ 3.2 Stage 2: Synthetic Observations
In Stage 2, the agent’s behavior is inferred on pre-generated synthetic observations. The scoring pipeline from Stage 1 is repeated for each of these synthetic observations. Stage 2 scores correspond to a range of plausible futures. We propose to weight their contributions towards a final combined score based on the proximity of Stage 2 start points to the Stage 1 endpoint. This prioritizes futures that are more likely. We show some examples of such generated scenes in Fig. 2. In the following, we provide details regarding the scenario pre-generation, scoring, and score aggregation processes.
@@
Score Aggregation. We score each synthetic observation using the EPDMS from Eq. 1 to obtain Stage 2 scores $\{s_{2}^{i}\}$. To compute the final score $s_{\text{combined}}$, we define two aggregation functions:
$s_{\text{combined}}=\mathcal{A}_{\text{1}}(s_{1},s_{2}),\quad\text{where}\ s_{2}=\mathcal{A}_{\text{2}}(\{s_{2}^{i}\},\{x^{i}\},\hat{x})$ (3)
$\mathcal{A}_{\text{1}}$ fuses the Stage 1 score $s_{1}$ with an aggregated Stage 2 result $s_{2}$. $\mathcal{A}_{\text{2}}$, in turn, aggregates $\{s_{2}^{i}\}$ based on their initial positions $\{x^{i}\}$, which denote the start points of the $i$-th Stage 2 scenario. $\hat{x}$ is the ego agent’s endpoint reached at the end of the Stage 1 simulation. In our experiments, we conduct an empirical study on different aggregation functions. Based on our findings, we instantiate $\mathcal{A}_{\text{1}}$ as a simple product, and $\mathcal{A}_{\text{2}}$ as a Gaussian-weighted average with kernel variance $\sigma^{2}$:
@@ 4 Results
In this section, we study three key questions to examine our proposed idea of pseudo-simulation.
@@ 4.1 How well-aligned is pseudo-simulation with closed-loop evaluation?
Results. First, we create a scatter plot comparing the 8-second closed-loop scores from the nuPlan simulator against the results of our 2$\times$4 second pseudo-simulation, which aims to approximate these closed-loop scores. As shown in Fig. 3 (a), pseudo-simulation exhibits strong correlation with closed-loop results across a broad range of planners, particularly among learned planners.
@@
Finally, we examine the effect of limiting the number of synthetic views in Stage 2.  Fig. 3 (e) reports correlation values when using 100%, 50%, and 25% of our available synthetic viewpoints. At 100% density, each scenario contains 12 synthetic observations on average in Stage 2 for each real observation in Stage 1, resulting in 13 planner inferences per scenario. In comparison, closed-loop simulation in nuPlan requires 80 planner inferences per scenario, corresponding to an 8-second rollout at 10Hz. This is 6$\times$ higher than pseudo-simulation. While subsampling reduces the number of synthetic views, the correlation to closed-loop scores remains strong. Even when using only 25% density, i.e., approximately three Stage 2 observations per scene, the correlation remains above 0.85. This indicates that pseudo-simulation maintains reliability even with reduced observation coverage.
@@ 4.2 What new challenges and insights does our leaderboard provide?
We select four baseline planners with varying input modalities. We aim to compare simple heuristics, state-of-the-art vision-based planners, and privileged systems with access to ground-truth information. As heuristics, we include the Constant Velocity (CV) and ego-history conditioned MLP from [8]. Representing vision-based methods, we evaluate Latent TransFuser (LTF) [11], the strongest publicly available image-only planner for nuPlan. Finally, PDM-Closed (PDM-C) [12], the best privileged planner on nuPlan, serves as our privileged system.
Results. Table 2 presents the detailed subscores for each planner, broken down by Stage 1 (original observations) and Stage 2 (synthetic observations). In terms of overall performance, PDM-Closed achieves the highest combined EPDMS of 51.3, followed by LTF (23.1), MLP (12.7), and CV (10.9). Comparing performance across stages, we observe a general drop in subscores from Stage 1 to Stage 2 for all methods (particularly for LTF), suggesting their sensitivity to the distribution shifts introduced in Stage 2. Notably, while PDM-Closed excels in most metrics, it exhibits lower performance in comfort metrics such as HC and EC. Our evaluation on navhard reveals this specific failure mode of PDM-Closed, highlighting a trade-off that was overlooked in prior benchmarks. We will host navhard as a public leaderboard for the community.
@@ 5 Conclusion
Correlation with Real-World Deployment. Our current validation focuses on establishing correlation with established simulation benchmarks. We do not yet demonstrate or claim direct correlation with performance metrics from real-world vehicle deployment. Bridging this gap between simulation-based evaluation and predicting real-world outcomes remains an important direction for future investigation.
@@
Rendering Fidelity and Evaluation. Despite achieving excellent quantitative results on rendering fidelity (LPIPS) and downstream task performance (mIoU, EPDMS), some visual artifacts may persist in the generated synthetic views. Our evaluation primarily focuses on algorithmic metrics. Future work may also benefit from incorporating human perceptual studies to gain a more comprehensive understanding of perceived realism and the potential impact of any remaining artifacts. Furthermore, combining neural rendering techniques like ours with state-of-the-art generative diffusion models might offer possibilities for enhancing rendering quality [58, 59, 60].

### v2

@@ 1 Introduction
However, accurate simulation remains a significant challenge, particularly for vision-based end-to-end AV systems. Real-world driving is visually complex and behaviorally diverse, making it difficult to replicate in simulation. Most existing platforms are manually constructed by 3D artists and engineers [3, 5]. This limits their realism and the diversity across scenes. Moreover, simulation-based evaluation is a computationally intensive and inherently sequential process. It often relies on large amounts of correlated evaluation frame sequences due to the high frequency of simulation required to ensure fidelity (usually 10Hz or higher).
Open-loop evaluation, on the other hand, measures planning performance by comparing predicted trajectories to expert demonstrations in pre-recorded datasets. Each observation for evaluation includes sensor inputs, a goal location, and the future trajectory executed by a human expert driver. The AV predicts a fixed-horizon trajectory conditioned on the inputs, which is then scored against the expert using either displacement errors or metrics derived from ground-truth (GT) environment annotations, such as lane compliance or estimated collisions [6, 7, 8]. This approach operates entirely on real sensor data and avoids the complexities of interactive simulation, making it scalable and straightforward to apply over large datasets. However, it evaluates behavior only under expert-aligned conditions and does not account for distribution shifts. In deployment, the AV deviates from the demonstrated path, and open-loop protocols do not test its ability to recover from such drift.
@@
Property Open-Loop Closed-Loop Pseudo-Sim
Real data Yes No Yes
Per-scene pre-processing No Yes Yes
Synthetic data rendering No Online Pre-rendered
Evaluation Parallel Sequential Parallel
Compounding errors No Yes Yes
Causal confusion No Yes Yes
In summary, pseudo-simulation combines real and pre-rendered synthetic data, enabling scalable, parallel evaluation.
[…]
@@ 2 Related Work
Closed-Loop Benchmarking. Graphics-based simulators support closed-loop evaluation, but are computationally expensive and introduce domain gaps in sensor fidelity [3, 5]. To improve scalability, data-driven planning simulators leverage recorded traffic data [4, 24, 25, 26]. However, these systems operate at the trajectory level and do not support sensor-based agents. Several works attempt to bridge this gap through data-driven sensor simulation, generating synthetic views from real-world logs. Early systems simulate ego-vehicle deviations via image-based rendering [27, 28, 29]. More recent methods explore neural rendering [30, 31, 32, 33, 34, 35]. However, they face challenges related to photorealism and runtime efficiency. As such, there is no widely established neural rendering based AV benchmark yet, which we aim to address in this work.
Open-Loop Benchmarking. Open-loop evaluation typically measures planning quality via displacement errors between predicted and expert trajectories [36]. These metrics are simple to compute, but often correlate poorly with real-world performance and tend to favor trivial or history-based baselines [7, 12, 37, 38]. Furthermore, benchmarks adopting the nuScenes dataset [39] exhibit inconsistencies in implementations for metrics such as ADE (Average Displacement Error) and collisions [6, 7], and overrepresent low-complexity scenes such as straight driving [7]. The closest work to ours, NAVSIM v1 [8], offers a more structured framework for open-loop benchmarking. The agent-under-test predicts a fixed-horizon trajectory from real sensor inputs, while other actors replay their recorded motion. This setup supports scalable evaluation and enables simulation-based metrics such as progress and collision rates. However, unlike pseudo-simulation, NAVSIM v1 remains limited to open-loop evaluation from expert-aligned initial observations and does not account for compounding errors or causal confusion [40].
@@ 3 Pseudo-Simulation
We consider a planning task where evaluation proceeds in two stages (Fig. 1). In both stages, an AV (also called planner/ego agent) generates a 4-second trajectory based on sensor inputs and a driving command [8]. The inputs include multi-view camera images and ego status features such as the velocity and motion history. The driving command specifies the intended maneuver in case of ambiguity, e.g., at intersections, and is provided as a discrete label: left, straight, or right. The ego agent outputs a trajectory (i.e., a sequence of desired future waypoints) in its local coordinate frame.
@@ 3.1 Stage 1: Initial Observations
Here, $\mathcal{M}_{\text{pen}}=\{\text{NC},\text{DAC},\text{DDC},\text{TLC}\}$ and $\mathcal{M}_{\text{avg}}=\{\text{TTC},\text{EP},\text{HC},\text{LK},\text{EC}\}$ (Table 1). Unlike prior work [44], to prevent penalizing contextually justified maneuvers, we introduce a novel filtering mechanism ($\text{filter}_{m}$) for the EPDMS. If a rule violation is also committed by the human expert driver in the same scene, the penalty is ignored. This avoids penalizing infractions due to label noise or valid behaviors, such as briefly entering the opposite lane to bypass a static obstacle.
@@ 3.2 Stage 2: Synthetic Observations
In Stage 2, the agent’s behavior is inferred on pre-generated synthetic observations. The scoring pipeline from Stage 1 is repeated for each of these synthetic observations. Stage 2 scores correspond to a range of plausible futures. We propose to weight their contributions towards a final combined score based on the proximity of Stage 2 start points to the Stage 1 endpoint. This prioritizes futures that are more likely. We show some examples of such generated scenes in Fig. 2. In the following, we provide details regarding the scenario pre-generation, scoring, and score aggregation processes. Note that we choose to create synthetic observations after unrolling for 4 seconds, instead of directly at the Stage 1 start point, since (1) this allows background traffic to react to the updated ego state, and (2) it provides a physically plausible history trajectory, which is a required planner input.
@@
[…]
@@ 4.1 How well-aligned is pseudo-simulation with closed-loop evaluation?
[…]
@@
[…]
@@ 4.2 What new challenges and insights does our leaderboard provide?
We select four baseline planners with varying input modalities. These include the Constant Velocity (CV) and Ego-history MLP from [8], Latent TransFuser (LTF) [11], the strongest public image-only planner for nuPlan, and, PDM-Closed (PDM-C) [12], the best privileged planner on nuPlan.
Results. Table 2 presents the detailed subscores for each planner, broken down by Stage 1 (original observations) and Stage 2 (synthetic observations). In terms of overall performance, PDM-Closed achieves the highest combined EPDMS of 51.3, followed by LTF (23.1), MLP (12.7), and CV (10.9). Comparing performance across stages, we observe a general drop in subscores from Stage 1 to Stage 2 for all methods (particularly for LTF), suggesting their sensitivity to the distribution shifts introduced in Stage 2. Notably, while PDM-Closed excels in most metrics, it exhibits lower performance in comfort metrics such as HC and EC. Our evaluation on navhard reveals this specific failure mode of PDM-Closed, highlighting a trade-off that was overlooked in prior benchmarks. We host navhard as a public leaderboard.
@@ 5 Conclusion
Correlation with Real-World Deployment. Our current validation focuses on establishing correlation with established simulation benchmarks. We do not yet demonstrate or claim direct correlation with performance metrics from real-world vehicle deployment. Bridging this gap between simulation-based evaluation and predicting real-world outcomes remains an important direction for future investigation. Rather than replacing real-world validation, frameworks to augment real-world evaluations with simulation can be applied more effectively with our work [56]
@@
Rendering Fidelity and Evaluation. Despite achieving excellent quantitative results on rendering fidelity (LPIPS) and downstream task performance (mIoU, EPDMS), some visual artifacts may persist in the generated synthetic views. Our evaluation primarily focuses on algorithmic metrics. Future work may also benefit from incorporating human perceptual studies to gain a more comprehensive understanding of perceived realism and the potential impact of any remaining artifacts. Furthermore, combining neural rendering techniques like ours with state-of-the-art generative diffusion models might offer possibilities for enhancing rendering quality [59, 60, 61, 62].
@@
Human-flag Filtering. Our filtering strategy disregards rule violations also committed by human experts. While this helps reduce false positives, it could also risk overlooking important failure and edge cases, since human driving is not always a gold standard for safety. Future work could further refine the human-flag filtering and explore this trade-off to ensure more reliable evaluation.
Metric Design Choices. We choose multiplicative aggregation because most sub-scores are binary-valued, and multiplication captures compounding failures, e.g., a collision should significantly impact the final score. Our Gaussian weighting is selected for its strong empirical performance with minimal assumptions. Exploring more principled formulations for aggregation and weighting remains an interesting future direction.